A chemist has two alloys, one of which is 5% gold and 15% lead and the other which is 40% gold and 50% lead. How many grams of each of the two alloys should be used to make an alloy that contains 42.5 g of gold and 92.5 g of lead?
Answer by lwsshak3(11628)

A chemist has two alloys, one of which is 5% gold and 15% lead and the other which is 40% gold and 50% lead. How many grams of each of the two alloys should be used to make an alloy that contains 42.5 g of gold and 92.5 g of lead?
***
let x=amt of 40% gold and 50% lead alloy to be used
let y=amt of 5% gold and 15% lead alloy to be used
..
50%x+15%y=92.5
40%x+5%y=42.5
..
50%x+15%y=92.5
120%x+15%y=127.5
subtract
-70%x=35
x=35/0.7=50
5%y=42.5-40%*50=42.5-20=22.5
y=22.5/.05=450
..
amt of 40% gold and 50% lead alloy to be used=50 g
let y=amt of 5% gold and 15% lead alloy to be used=450 g